# Classification of Hoyle State Decay Branches in Active Target Time Projection Chamber using Neural Network

### v1

@@ 3 Performance Analysis
In this section, the evaluation of the training and validation, followed by classification performance of the model, have been discussed. The effect of variation in readout segmentation of the SAT-TPC has also been studied.
@@ 3.1 Training & Validation Accuracy
To evaluate the training of the present CNN model using regularization technique, the accuracy and loss curves of training and validation have been produced, as shown in figure 8. The curves show that the performance of the model has improved as the number of epoch has increased and stabilized after 10.
@@ 3.2 Classification Performance
We have evaluated the performance of the CNN model using three key metrics: precision, recall, and F1 score. These metrics have provided more detailed insights into its performance and errors than accuracy alone. For the track classification, we have focused on the ability of the model to identify direct and sequential decay events and the backgrounds accurately. The precision is a measure of the accuracy of a model in identifying the positive instances which is calculated as the ratio of correctly predicted positive observations to the total predicted positives. In simpler terms, it tells us how many of the items that the model identifies as positive are actually positive. Recall, also known as sensitivity, measures the ability of the model to identify all relevant instances within a dataset. It is calculated as the ratio of correctly predicted positive observations to all observations in the actual class. Essentially, it indicates how many of the actual positives the model is able to capture. The F1 score is the harmonic mean of precision and recall. It provides a single metric that balances both the precision and recall of a model, especially useful when the classes are imbalanced. The F1 score combines the strengths of precision and recall, making it a comprehensive metric to evaluate the performance of a model. It is useful particularly in situations where the cost of false positives and false negatives is significant. The support column indicates the number of true instances (actual samples) of each event class in the dataset, used to evaluate the model. The performance metrics of the present model are given in table 2. These have been determined with a readout segmentation of 2 mm $\times$ 2 mm.
@@
The two metrics, precision and recall, are best understood by producing a confusion matrix as shown in figure 9 which facilitates assessment of the performance and errors of the CNN model. It is evident from the matrix that the model has exhibited strong performance in classifying the scattering events $\alpha$ + 40Ar, $\alpha$ + 12C, and $\alpha$ + 16O, with high accuracy. The matrix reveals the misclassifications between the direct and sequential decays which indicates that these classes might require more distinct training data. In order to test this concept, the readout segmentation has been varied to study its effect on the performance of the model (discussed in the next sub-section).
@@ 3.3 Effect of Readout Segmentation
The scheme of the anode plane segmentation of the SAT-TPC in the simulation model has been varied to 1 mm $\times$ 1 mm and 3 mm $\times$ 3 mm pixels. The calculation has been carried out with the same class of events, discussed in the previous section. The classification performance for the two segmentation schemes in terms of the three metrics is given in table 3. It shows that the reduction in the size of the readout pixel from 3 mm $\times$ 3 mm to 1 mm $\times$ 1 mm has led to improved classification accuracy, particularly for decay events. This suggests that finer segmentation enhances spatial resolution, allowing the CNN to distinguish subtle differences in track morphology.
Table 3: Classification report for segmentation schemes 1 mm $\times$ 1 mm and 3 mm $\times$ 3 mm
@@
1 mm $\times$ 1 mm direct decay 0.62 0.67 0.64 200
@@
3 mm $\times$ 3 mm direct decay 0.90 0.82 0.86 200
@@ 4 Summary & Conclusion
The classification of Hoyle state decay events and the background elastic scattering events forms the central theme of the present study, showcasing the application of supervised machine learning, particularly CNN, for accurate segregation of the relevant class of events. The work has highlighted the challenges of distinguishing between direct decay, sequential decay, and the background scattering events $\alpha$ + 40Ar, $\alpha$ + 12C, and $\alpha$ + 16O in the context of using 30 MeV $\alpha$-beam and a prototype SAT-TPC, filled with active gas target Ar + CO2. We have performed Monte Carlo simulations to study the nuclear interactions leading to formation of the Hoyle state, following excitation through inelastic scattering of $\alpha$ from the active target nucleus 12C and the relevant elastic scattering of $\alpha$ from other target nuclei 40Ar, 12C and 16O. The decay kinematics have been analyzed in a boosted frame, allowing the extraction of energy and angular distributions of the emitted $\alpha$-particles. The same has been done for the elastic scattering events to study the energy and angular distributions of the scattered products. The primary ionization by the beam and the scattering and decay products has been simulated in Geant4 using a low-energy physics list that has been used to reconstruct the event tracks in the SAT-TPC active volume, convoluted with diffusion parameter and projected on two-dimensional readout with segmentation 2 mm $\times$ 2 mm. Using the VGG-16 CNN architecture, the model has classified the events on the basis of the track morphology images. Regularization technique, such as dropout, has been employed to prevent overfitting and enhance the robustness of the model, achieving high precision, recall, and F1 scores across all categories.
@@
Despite strong classification performance, the study has identified specific trends in misclassifications, particularly among the decay events. The classification performance of the model for the decay events has been found to improve with the finer readout segmentation as it has reduced the misclassification rate of the decay events. These results have proved the efficiency of the model in distinguishing Hoyle state decays from the background events while highlighting areas for further optimization, such as refining feature representations and detector configurations to enhance separation between closely related scattering categories.
@@
This work has demonstrated the utilization of CNN in analyzing experimental data of low-energy nuclear physics experiments performed with SAT-TPC, providing a framework for advancing robust event classification in nuclear and astrophysical research. Future work will explore advanced architectures, such as ResNet, and transformer-based models, to further enhance classification performance. In addition, real-time processing capabilities will be investigated to enable integration with experimental data acquisition systems.
@@ Acknowledgments
The authors acknowledge the support provided by their institute. They are thankful to Dr. Tapan Kumar Rana and Dr. Santu Manna of the Variable Energy Cyclotron Center, and Prof. Satyaki Bhattacharya of Saha Institute of Nuclear Physics Kolkata, India, for valuable discussion and suggestions.

### v2

@@ 3 Performance Analysis
In this section, the evaluation of the training and validation, followed by classification performance of the model, are discussed. The effect of variation in readout segmentation of the SAT-TPC has also been studied.
@@ 3.1 Training & Validation Accuracy
To evaluate the training of the present CNN model using a regularization technique, the accuracy and loss curves of training and validation have been produced, as shown in figure 8. The curves show that the performance of the model has improved as the number of epochs has increased and stabilized after 10.
@@ 3.2 Classification Performance
We have evaluated the performance of the CNN model using three key metrics: precision, recall, and F1 score. These metrics have provided more detailed insights into its performance and errors than accuracy alone. For the track classification, we have focused on the ability of the model to identify direct and sequential decay events and the backgrounds accurately. The precision is a measure of the accuracy of a model in identifying the positive instances which is calculated as the ratio of correctly predicted positive observations to the total predicted positives. In simpler terms, it tells us how many of the items that the model identifies as positive are actually positive. Recall, also known as sensitivity, measures the ability of the model to identify all relevant instances within a dataset. It is calculated as the ratio of correctly predicted positive observations to all observations in the actual class. Essentially, it indicates how many of the actual positives the model is able to capture. The F1 score is the harmonic mean of precision and recall. It provides a single metric that balances both the precision and recall of a model, especially useful when the classes are imbalanced. The F1 score combines the strengths of precision and recall, making it a comprehensive metric to evaluate the performance of a model. It is useful particularly in situations where the cost of false positives and false negatives is significant. The support column indicates the number of true instances (actual samples) of each event class in the dataset used to evaluate the model. The performance metrics of the present model are given in table 2. These have been determined with a readout segmentation of 2 mm $\times$ 2 mm.
@@
The two metrics, precision and recall, are best understood by producing a confusion matrix, as shown in figure 9 which facilitates assessment of the performance and errors of the CNN model. It is evident from the matrix that the model has exhibited strong performance in classifying the scattering events $\alpha$ + 40Ar, $\alpha$ + 12C, and $\alpha$ + 16O, with high accuracy. The matrix reveals the misclassifications between the direct and sequential decays which indicates that these classes might require more distinct training data. In order to test this concept, the readout segmentation has been varied to study its effect on the performance of the model (discussed in the next sub-section).
@@ 3.3 Effect of Readout Segmentation
The scheme of the anode plane segmentation of the SAT-TPC in the simulation model has been varied to 1 mm $\times$ 1 mm and 3 mm $\times$ 3 mm pixels. The calculation has been carried out with the same class of events discussed in the previous section. The classification performance for the two segmentation schemes in terms of the three metrics is given in table 3. It shows that the reduction in the size of the readout pixel from 3 mm $\times$ 3 mm to 1 mm $\times$ 1 mm has led to improved classification accuracy, particularly for decay events. This suggests that finer segmentation enhances spatial resolution, allowing the CNN to distinguish subtle differences in track morphology.
Table 3: Classification report for segmentation schemes 3 mm $\times$ 3 mm and 1 mm $\times$ 1 mm
@@
3 mm $\times$ 3 mm direct decay 0.62 0.67 0.64 200
@@
1 mm $\times$ 1 mm direct decay 0.90 0.82 0.86 200
@@ 4 Summary & Conclusion
The classification of Hoyle state decay events and background elastic scattering events forms the central theme of the present study, showcasing the application of a supervised machine learning technique, specifically a CNN, for accurate segregation of the relevant class of events. The work has highlighted the challenges of distinguishing between direct decay, sequential decay, and background scattering events $\alpha$ + 40Ar, $\alpha$ + 12C, and $\alpha$ + 16O in the context of using 30 MeV $\alpha$-beam and a prototype SAT-TPC, filled with active gas target Ar + CO2. We have performed Monte Carlo simulations to study the nuclear interactions leading to formation of the Hoyle state, following excitation through inelastic scattering of $\alpha$ from the active target nucleus 12C and the relevant elastic scattering of $\alpha$ from the target nuclei 40Ar, 12C and 16O. The decay kinematics have been analyzed in a boosted frame, allowing for the extraction of energy and angular distributions of the emitted $\alpha$-particles. The same has been done for the elastic scattering events to study the energy and angular distributions of the scattered products. The primary ionization by the beam and the scattering and decay products has been simulated in Geant4 using a low-energy physics list that has been used to reconstruct the event tracks in the SAT-TPC active volume, convoluted with a diffusion parameter, and projected on 2D readout with segmentation of 2 mm $\times$ 2 mm. Using the VGG-16 CNN architecture, the model has classified the events on the basis of the track morphology images. Dropout was employed as a regularization technique to prevent overfitting and enhance the robustness of the model, achieving high precision, recall, and F1 scores across all categories.
@@
Despite strong classification performance, the study has identified specific trends in misclassifications, particularly among the decay events. The classification performance of the model for the decay events has been found to improve with the finer readout segmentation as it has reduced the misclassification rate of the decay events. These results have proved the efficiency of the model in distinguishing Hoyle state decays from the background events while highlighting areas for further optimization, such as refining feature representations and detector configurations to enhance separation between closely related scattering categories. In the future, the size of the dataset will be increased well above the current choice of 1000 events/class to enhance training diversity and classification performance, especially for the more ambiguous decay modes. Additionally, since the CNN is trained on Monte Carlo data, further investigation will be needed to evaluate its robustness and performance on real detector data, including domain adaptation or transfer learning approaches.
@@
This work has demonstrated the utilization of a CNN in analyzing experimental data of low-energy nuclear physics experiments performed with the SAT-TPC, providing a framework for advancing robust event classification in nuclear and astrophysical research. Future work will explore advanced architectures, such as ResNet, and transformer-based models, to further enhance classification performance. In addition, real-time processing capabilities will be investigated to enable integration with experimental data acquisition systems.
@@ Acknowledgments
The authors acknowledge the support provided by their institute. They are thankful to Dr. Tapan Kumar Rana and Dr. Santu Manna of the Variable Energy Cyclotron Center, Kolkata and Prof. Satyaki Bhattacharya of Saha Institute of Nuclear Physics, Kolkata for valuable discussions and suggestions.